OpenStudy (anonymous):

Lizbeth deposited $6200 into a savings account for which interest is compounded monthly at a rate of 2.37%. How much interest will she earn after 8 years?

OpenStudy (paki):

I=PRT/100 I= 1140$

OpenStudy (ranga):

Here is the compound interest formula: \[\Large A = P(1 + \frac{ r }{ n })^{nt}\]A = Amount at maturity P = Principal Amount r = Annual interest rate in decimal n = compounding period (compounded how many times a year) t = years invested Calculate A. Subtract P to find the interest earned.

OpenStudy (ranga):

Here, P = 6200; r = 0.0237; n = 12; t = 8. Plug the numbers into the compound interest formula, calculate A. Interest earned = A - P = A - 6200 = ?

OpenStudy (anonymous):

i got an answer that i know isnt right . 590175.52

OpenStudy (ranga):

Make proper use of parenthesis in your calculator. Or to avoid complicated entry simplify each number first.

OpenStudy (ranga):

A = 6200(1 + 0.0237/12)^(12*8) = 6200(1.001975)^96 = ?

OpenStudy (anonymous):

after that the problem would look like 6212.245^96 right ?

OpenStudy (ranga):

Be careful when you are using the calculator. Tell me what you are entering in each step.

OpenStudy (ranga):

Enter: 1.001975 x^y 96 = x 6200 =